OpenStudy (anonymous):

pat purchased a 23,000 car in a city with 8% sales tax, another city has 9.25 sales tax... how much did he save?

OpenStudy (anonymous):

\[$23000 * 0.08=$1840\] \[$23000 * 0.0925 = 2127.5\] \[2127.5-1840=$287.5\]

OpenStudy (anonymous):

thanks :3